# Attention-based Shape-Deformation Networks for Artifact-Free Geometry Reconstruction of Lumbar Spine from MR Images

### v1

@@ 4.2.2 UNet-Disp Model
This model was originally proposed by Daniel et al. [24] for creating simulation-ready heart valve geometries from 3D CT images. Compared to the UNet-GCN model, the UNet-Disp model also uses a UNet but in a different way: the decoder of the UNet predicts a displacement field that is used to deform a template. For our application, we followed the original design as close as possible, and made necessary adjustment by changing the 3D UNet to a 2D UNet that outputs a 2D displacement field. The UNet-Disp model does not use any attention mechanisms.
@@ 4.2.3 UNet-MLP Model
This model is a simplification of our UNet-DeformSA model by replacing shape self-attention layers with a simple MLP.
@@ 4.4 Metrics and Results of Geometry Reconstruction
To measure the accuracy of a reconstructed lumbar object (a vertebra or a disc), we assessed two key metrics: (1) Averaged Point-to-Point Distance (APPD) between a predicted shape and the corresponding ground-truth shape and (2) Dice Similarity Coefficient (Dice), measuring the overlap between areas enclosed by predicted and ground-truth shapes. For each metric, mean, standard deviation, the worse case, $95$th or $5$th percentile among the samples in the test set are calculated.
@@
Figure 8: Outputs of the models on two challenging cases. Different colors indicate different lumbar components. On the upper-right corner, the white number shows the average point-to-point error (mm), and the yellow number shows the maximum point-to-point error (mm).
@@ 6 Conclusion
We proposed two novel attention-based neural networks, UNet-DeformSA and TransDeformer, to automatically reconstruct lumbar spine geometries with mesh correspondence from 2D MR images. The reconstructed geometries are highly accurate and free of artifacts. In addition, we proposed a shape error estimation network based on TransDeformer, which facilitates quality control. Thus, we have provided a complete solution for fast and accurate measurement of the key medical parameters of lumbar spinal components from MR images.

### v2

@@ 4.2.2 UNet-Disp Model
This model was originally proposed by Pak et al. [24] for creating simulation-ready heart valve geometries from 3D CT images. Compared to the UNet-GCN model, the UNet-Disp model also uses a UNet but in a different way: the decoder of the UNet predicts a displacement field that is used to deform a template. For our application, we followed the original design as close as possible, and made necessary adjustment by changing the 3D UNet to a 2D UNet that outputs a 2D displacement field. The UNet-Disp model does not use any attention mechanisms.
@@ 4.2.3 UNet-MLP Model
This model is a simplification of our UNet-DeformSA by replacing shape self-attention layers with a simple MLP in each geometry-deformation module.
@@ 4.4 Metrics and Results of Geometry Reconstruction
To measure the accuracy of a reconstructed lumbar object (a vertebra or a disc), we assessed two key metrics: (1) Averaged Point-to-Point Distance (APPD) between a predicted shape and the corresponding ground-truth shape and (2) Dice Similarity Coefficient (Dice), measuring the overlap between areas enclosed by predicted and ground-truth shapes. For each metric, mean, standard deviation (std), the worse case, the $95$th or $5$th percentile among the test samples are calculated.
@@
Figure 9: Outputs of the models on two challenging cases. Different colors indicate different lumbar components. On the upper-right corner, the white number shows the average point-to-point distance error (mm), and the yellow number shows the maximum point-to-point distance error (mm).
@@ 6 Conclusion
We proposed two novel attention-based neural networks, UNet-DeformSA and TransDeformer, to automatically reconstruct lumbar spine geometries with mesh correspondence from 2D MR images. The reconstructed geometries are highly accurate and free of artifacts. In addition, we proposed a shape error estimation network based on TransDeformer, which facilitates quality control. Thus, we have provided a complete solution for fast and accurate measurement of the key medical parameters of lumbar spine components from MR images.